OpenStudy (anonymous):

Prove using mathematical induction that 1^3+2^3+...+n^3=(n^2(n+1)^2)/4

OpenStudy (freckles):

You must do it for base case first.

OpenStudy (freckles):

Then after that: \[\text{ Assume } 1^3+2^3+ \cdots +k^3=\frac{k^2(k+1)^2}{4} \text{ holds for some integer } \\ k>0\] \[\text{ And show } 1^3+2^3+\cdots +(k+1)^3 =\frac{(k+1)^2(k+2)^2}{4} \text{ holds }\]

OpenStudy (freckles):

hint: \[1^3+2^3+\cdots +(k+1)^3 \\ =1^3+2^3+ \cdots +k^3 +(k+1)^3 \\ =\frac{k^2(k+1)^2}{4}+(k+1)^3 \]

OpenStudy (freckles):

you must use algebra to show that equals (k+1)^2(k+2)^2/4

OpenStudy (freckles):

david do not make it harder than it needs to be do not multiply all of that out

OpenStudy (freckles):

Let me know when and if you need a hint

OpenStudy (anonymous):

I just don't understand when I reached Part Three of Induction Proofs "Prove true for k+1"

OpenStudy (freckles):

You are saying you don't understand what it means?

OpenStudy (anonymous):

I do. I just don't know how to get the final answer.

OpenStudy (freckles):

Oh first step take your second term and multiply it by 4/4

OpenStudy (freckles):

\[=\frac{k^2(k+1)^2}{4}+\frac{4(k+1)^3}{4}\]

OpenStudy (freckles):

Doing this will allow us to write as one fraction

OpenStudy (freckles):

\[=\frac{k^2(k+1)^2+4(k+1)^3}{4}\]

OpenStudy (anonymous):

See. That is what I don't understand.

OpenStudy (freckles):

the combining fractions part?

OpenStudy (anonymous):

How about the first part \[1^3+2^3+...+n^3=\]

OpenStudy (anonymous):

How does this work out?

OpenStudy (freckles):

Are we talking about the base case? showing it is true for n=1?

OpenStudy (freckles):

Step 1: Show it is true for the base case. The base case for this one is showing it is true for n=1: \[1^3 =1 =\frac{1^2(1+1)^2}{4} \text{ holds } \]

OpenStudy (anonymous):

Base case? I don't think so. What I mean is how does it expand? For example:\[1+2+3+...+k= \] becomes \[1+2+3+k+ k+1\] This is what I mean. How does \[1^3+2^3+...+n^3 \] become "expanded"??

OpenStudy (freckles):

\[\text{ Assume } 1^3+2^3+ \cdots +k^3=\frac{k^2(k+1)^2}{4} \text{ holds for some integer } \\ k>0 \text{ And show } 1^3+2^3+\cdots +(k+1)^3 =\frac{(k+1)^2(k+2)^2}{4} \text{ holds }\] That thing I assumed is true (first line of this post) is an equation I found just from replacing n with k The thing we want to show using the thing we assume is true is an equation I found from just replacing the n's with (k+1)'s

OpenStudy (freckles):

@amistre64 or @hero

OpenStudy (freckles):

I'm calling others here because I don't understand. I'm sorry.

OpenStudy (anonymous):

I don't understand either. I hate Mathematical Inductions.

OpenStudy (freckles):

Like I mean I don't understand your question is all.

OpenStudy (amistre64):

proof by induction is all about form .... since know a certain form is true we must manipulate the k+1th element into the 'true form'

OpenStudy (anonymous):

Like this: \[1+2+3+...=\frac{ n(n+1) }{ 2 }\] \[1+2+3+...+k+k+1= \frac{ (k+1)(k+1+1) }{ 2 }\]\[\frac{ k(k+1) }{ 2} + k+1 = \frac{ (k+1)(k+2) }{ 2 }\] \[\frac{ k(k+1)+2(k+1) }{ 2 } = \frac{ (k+1)(k+2) }{ 2}\] See how it's "expanding" after the first line Freckles???

OpenStudy (amistre64):

after the basis step, which proves its true for a specific value, we want to make it general again ... let n=k 1^3+2^3+...+k^3 = (k^2(k+1)^2)/4 <= it is this FORM that we know is good if we let the left side be named, say P(k) then P(k) = (k^2(k+1)^2)/4 add then k+1th term to each side .. P(k) + (k+1)^3 = (k^2(k+1)^2)/4 + (k+1) now the leftside is just the P(k+1) term, and we can focus on the right side manipulating it into the 'good' form to show that k has turned into k+1

OpenStudy (amistre64):

forgot my ^3 on the right side :)

OpenStudy (anonymous):

Ok. I kind of starting to get it.

OpenStudy (freckles):

maybe if I do an example for you or something... Let's show \[1^2+2^2+3^2 + \cdots + n^2=\frac{n(n+1)(2n+1)}{6} \] The first case is showing it is true for the base case. The base case being for n=1 in this case So we have \[1^2 \text{ on left hand side } \text{ and } \frac{1(1+1)(2(1)+1)}{6}=\frac{1(2)(3)}{6} \\ \text{ both sides are equal to 1 } \text{ this shows the base case }\] Now we have an if then statement \[\text{ If } 1^2+2^2+\cdots + k^2=\frac{k(k+1)(2k+1)}{6} \\ \text{ for some integer } k \ge 1 \text{ , then } \\ 1^2+2^2+\cdots +(k+1)^2=\frac{(k+1)(k+2)(2(k+1)+1)}{6} \text{ holds }\] So in if then statements we get to suppose the if part and show the then part... \[1^2+2^2+\cdots +(k+1)^2 \\ \text{ we know k comes before (k+1) } \\ \text{ so we can write } \\ 1^2+2^2+ \cdots +k^2+(k+1)^2 \\ \text{ but by our induction hypothesis we can write } 1^2+2^2+ \cdots + k^2 \\ \text{ as } \frac{k(k+1)(2k+1)}{6} \\ \text{ so we have } 1^2+2^2 + \cdots +k^2+ (k+1)^2 \text{ can be written as } \\ \frac{k(k+1)(2k+1)}{6} +(k+1)^2 \\ \text{ Remember we are trying to show this is equal to the one fraction } \\ \frac{(k+1)(2k+1)(2k+3)}{6} \\ \text{ so let's look at } \\ \frac{k(k+1)(2k+1)}{6} +(k+1)^2 \text{ and think ... } \\ \text{ well we need one fraction so ... let's combine fractions } \\ \frac{k(k+1)(2k+1)}{6} +\frac{6(k+1)^2}{6} \\ \frac{k(k+1)(2k+1)+6(k+1)^2}{6} \\ \text{ now do some factoring } \frac{(k+1)[k(2k+1)+6(k+1)}{6} \\ \frac{(k+1)[2k^2+k+6k+6]}{6} \\ \frac{(k+1)(2k^2+7k+6)}{6} \\ \frac{(k+1)(2k^2+3k+4k+6)}{6} \\ \frac{(k+1)[k(2k+3)+2(2k+3)]}{6} \\ \frac{(k+1)(2k+3)(k+2)}{6} \\ \text{ which is what we wanted to show } \\ \text{ so therefore we can conclude } \\ 1^2+2^2+\cdots +n^2=\frac{n(n+1)(2n+1)}{6} \text{ holds for all integer } n \ge 1. \]

OpenStudy (amistre64):

the thing is, if we knows it true for some k (we know its true for the base case) and can show by 'form' that the (k+1)th term fits the same format, then we have shown that if we simply replace k with k+1 that the setup remains consistent ... nothing is altered, and therefore we can say its true for the next term, and the next term, and the next ..

OpenStudy (amistre64):

term being defined as an integer that is

OpenStudy (freckles):

1+2+3+...+k doesn't become 1+2+3+...+k+(k+1)

OpenStudy (amistre64):

it does if you add the next term :) but other than that its immaterial what the left side looks like

OpenStudy (freckles):

saying 1+2+3+...+k becomes 1+2+3+...+k+(k+1) feels like someone is saying these are equal so I just didn't like the word becomes

OpenStudy (freckles):

The statement is P(n) and we want to show P(n) is true for all positive integer n So show P(1) is true assume P(k) is true show P(k+1) is true

OpenStudy (amistre64):

'becomes' has a connotation of change, of morphing, of changing by default ... not equality or sameness.

OpenStudy (amistre64):

tomato tomato tho lol

OpenStudy (freckles):

Ok ok like the spice girls say 2 becomes 1

OpenStudy (freckles):

2 becomes 1 by subtracting 1 from 2

OpenStudy (freckles):

or there some other ways 2 can become 1

OpenStudy (freckles):

But I don't know if I should rely on Spice Girls for my mathematical meanings

OpenStudy (amistre64):

she becomes pregnant infers something has changed :)

OpenStudy (amistre64):

david, any thoughts brewing?

OpenStudy (anonymous):

Well my math teacher taught us that k becomes k+k+1

OpenStudy (anonymous):

Why must Mathematical Inductions be taught anyways?? I don't see it being useful in life anyways.

OpenStudy (freckles):

Well I know it is useful in some computer science-y things and it definitely useful for proving things... It also exercises your algebra skills.

OpenStudy (freckles):

Or so I heard it is useful in computer science

OpenStudy (anonymous):

I'm more into Biology, that's why.

OpenStudy (freckles):

http://www.cs.sfu.ca/~bbart/225/induction1-4.html

OpenStudy (freckles):

This says why it is useful in computer science But I have no clue how it will help you in bio

OpenStudy (freckles):

Besides to help strengthen those critical thinking skills and algebra skills

OpenStudy (anonymous):

Exactly. This proves my point why I don't find this useful at all. I suck in math.

OpenStudy (freckles):

I think there is math in bio if I'm not mistaken

OpenStudy (freckles):

I definitely know statistics is important to every science

OpenStudy (anonymous):

There is math in bio but it's not the hard type of math, just some hard math now and then. But never have I seen mathematical inductions being used in bio